OpenStudy (anonymous):

Which inequality is equivalent to x – 5 ≤ 2x – 6? a. x ≤ 1 b. x ≥ -1 c. x ≥ 1 d. x ≥ 11 e. x ≤ 11 @iGreen help please?

OpenStudy (igreen):

\(x – 5 \le 2x – 6\) Add 5 to both sides: \(x \le 2x – 1\) Subtract 2x to both sides: \(-x \le -1\) Divide -1 to both sides. What's -1 / -1? @pigglet

OpenStudy (igreen):

If you're seeing those black question marks, click on a different post and click back here..or reload the page.

OpenStudy (anonymous):

1

OpenStudy (igreen):

Yep. So we have: \(x \ge 1\) Remember to switch the sign when dividing by a negative number.

OpenStudy (anonymous):

thanks